# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v2

@@ Abstract
We develop a rotation-invariant neural network that provides the global minimum-variance portfolio by jointly learning how to lag-transform historical returns and marginal volatilities and how to regularise the eigenvalues of large equity covariance matrices. This explicit mathematical mapping offers clear interpretability of each module’s role, so the model cannot be regarded as a pure black box. The architecture mirrors the analytical form of the global minimum-variance solution yet remains agnostic to dimension, so a single model can be calibrated on panels of a few hundred stocks and applied, without retraining, to one thousand US equities—a cross-sectional jump that indicates robust generalization capability. The loss function is the future short-term realized minimum portfolio variance and is optimized end-to-end on real daily returns. In out-of-sample tests from January 2000 to December 2024, the estimator delivers systematically lower realized volatility, smaller maximum drawdowns, and higher Sharpe ratios than the best analytical competitors, including state-of-the-art non-linear shrinkage, and these advantages persist across both short and long evaluation horizons despite the model’s training focus on minimizing short-term variance. Furthermore, although the model is trained end-to-end to produce an unconstrained (long-short) minimum-variance portfolio, we show that its learned covariance representation can be used in general optimizers under long-only constraints with virtually no loss in its performance advantage over competing estimators. These gains persist when the strategy is executed under a highly realistic implementation framework that models market orders at the auctions, empirical slippage, exchange fees, and financing charges for leverage, and they remain stable during episodes of acute market stress.
@@ 1 Introduction
Portfolio management has been fundamental to financial practice since the earliest securities markets emerged. Following Markowitz’s introduction of Modern Portfolio Theory (MPT) [1], academic and industry efforts have prioritized the construction of portfolios that optimize risk-adjusted returns. This framework employs historical return data and estimates of pairwise asset dependencies, typically captured by covariance matrices, to determine asset weightings.
@@
The literature on machine learning–based portfolio optimization is vast and rapidly expanding [4]. Existing studies can be grouped into three main categories: span predict-then-optimize pipelines, end-to-end differentiable frameworks, and reinforcement-learning agents. Overall, machine learning–based portfolio optimization is highly challenging with several limitations still to be overcome. In fact, some studies employ small or static asset universes that limit evidence of scalability and robustness. Benchmarks are in some cases infrequent or sub-optimal, making it difficult to compare performance improvements. Some designs inadvertently introduce data leakage through look-ahead bias or survivorship bias. These challenges constitute the principal obstacles to the practical deployment of machine learning in portfolio optimization.
The present study introduces an NN architecture that implements an end-to-end GMV portfolio optimizer by mapping each analytical step of the classical solution into a modular network. Historical returns for $n$ assets over $\Delta t$ time periods are first processed through a learnable lag-transformation module that applies distinct weights and soft-clipping parameters across look-back lags. The transformed returns then feed into a bidirectional LSTM–based correlation denoising module that adapts to the aspect ratio $q=n/\Delta t$, and into a parallel feed-forward network that estimates inverse marginal volatilities. All modules are calibrated jointly to minimize the Out-Of-Sample (OOS) portfolio variance in an end-to-end fashion. The correlation denoising module represents the greatest learning challenge, since its multivariate nature and the inherent symmetry of correlation matrices impose complex constraints on the network. If the architecture is correctly designed, calibration on datasets with $n$ in a narrow range should yield a parameter set that generalizes to different values of $n$. This expectation is borne out in our experiments. OOS tests on a universe of one thousand U.S. equities covering January 2000 through December 2024, using calibration on just a few hundred stocks, confirm that our architecture delivers lower realised variance and higher Sharpe ratios than state-of-the-art methods. When evaluated under realistic trading procedures and cost assumptions, it also produces consistently smaller drawdowns.
@@
A key aspect of our model is associated with the determination of the covariance matrix. When estimated from finite samples, empirical covariance matrices tend to be ill-conditioned and sample covariance presents estimation errors [3, 5, 6], often resulting in suboptimal portfolio allocations. A widely adopted remedy is to apply Rotation Invariant Estimators (RIEs), which adjust the sample eigenvalues while retaining the sample eigenvectors [7]. Within this framework, optimality is measured by the Mean Squared Error (MSE), equivalent to the Frobenius norm distance between the estimator and the true covariance matrix. The theoretical optimum is the so-called Oracle Estimator (OE), defined as the covariance matrix that minimizes this Frobenius distance. In practice, however, the OE cannot be realized, since it requires prior knowledge of the unknown true covariance.
@@
One of the earliest covariance regularization techniques is eigenvalue clipping. Laloux et al. [5] showed that the bulk of the eigenvalue spectrum of empirical financial covariance matrices closely follows the Marchenko–Pastur distribution for random noise. Eigenvalue clipping therefore replaces all sample eigenvalues that lie within this noise bulk by their average value. Later, Ledoit and Wolf [6] introduced the Linear Shrinkage (LS) estimator, which constructs a convex combination of the sample covariance matrix and a diagonal target matrix. This estimator remains well–defined when the number of assets $n$ exceeds the number of observations $\Delta t$. Its optimality in minimizing MSE holds strictly under the assumption that the true covariance matrix is drawn from an inverse–Wishart distribution [8]. In the last decade, Ledoit and collaborators introduced the NonLinear Shrinkage (NLS) estimator as an optimal method for covariance regularization [9]. Deriving the analytical form of the Ledoit–Péché (LP) estimator is challenging because it requires detailed knowledge of the sample eigenvector distribution [10]. The Quantized Eigenvalues Sampling Transform (QuEST) delivers the most accurate numerical implementation of the LP estimator by discretizing the limiting spectral distribution [11], but it incurs substantial computational cost. To alleviate this burden, the direct kernel estimator [12] and the inverse–Wishart regularization approach [7] offer closed-form approximations that closely mimic the LP solution. More recently, Ledoit and Wolf proposed the Quadratic-Inverse Shrinkage (QIS) estimator [13], which is considered the state-of-the-art technique in financial covariance filtering by combining computational efficiency with high estimation accuracy.
@@
Although the theoretical framework for covariance filtering is well established, its practical implementation remains challenging due to the pronounced non-stationarity of financial covariances. Market regimes evolve continuously under shifting economic conditions and investor behavior, so a covariance matrix estimated from past returns may poorly reflect current asset dependencies. Moreover, NLS theory assumes that sampling noise is the sole source of discrepancy between the sample and true covariance matrices and relies on the thermodynamic limit in which the number of assets $n$ and observations $\Delta t$ both tend to infinity with fixed aspect ratio $q=n/\Delta t$. In practice, extending $\Delta t$ to satisfy this limit only exacerbates non-stationarity, as longer historical windows encompass multiple structural shifts. Finally, the optimality of the LP estimator depends on the existence of finite twelfth-order moments of asset returns [10], yet empirical evidence indicates that equity returns exhibit heavy tails with tail exponents near four, rendering higher-order moments effectively unbounded. Methods based on $k$-fold Cross-Validation (CV) have been proposed to relax the stringent assumptions of nonlinear shrinkage theory [14, 15, 16, 8, 17]. These CV estimators achieve accuracy comparable to more complex schemes while permitting straightforward adaptations to specific data features. For example, Tan et al. [18] introduced an exponential decay factor into the return series within the CV framework. By applying exponential weights, one can maintain a long calibration window—approaching the thermodynamic limit—while emphasizing the most recent observations. This modification partially resolves the variance–bias trade-off caused by non-stationarity and structural shifts in financial covariances.
@@
An alternative to RIEs employs hierarchical clustering to filter the empirical correlation matrix [19, 20, 21]. Unlike RIEs, clustering-based filters modify both eigenvalues and eigenvectors by aggregating assets into empirically determined groups. This approach exploits the fact that financial assets are often organized into a clear hierarchy of sectors and sub-sectors [22]. Although it lacks full theoretical support, clustering filters have, in some instances, yielded superior out-of-sample portfolio performance compared with RIEs [23, 24, 20, 21, 25]. By enforcing the observed hierarchical relationships, these methods produce the most parsimonious covariance approximation that exactly reproduces the data’s cluster structure.
@@
The Engle Dynamical Conditional Covariance (DCC) model [26] represents a principled attempt to accommodate non-stationarity by modeling time-varying correlations through GARCH-type dynamics. It specifies separate equations for the conditional variances and then characterizes the evolving correlations of the standardized returns via a time-varying correlation matrix. In principle, this approach should adapt to shifting market regimes and capture persistence in asset comovements. In practice, however, DCC is computationally intensive, often suffers from convergence issues, and yields OOS portfolios whose risk–return profiles rarely surpass those obtained with simpler static filters [27]. In contrast, the Average Oracle (AO) method [28, 27] abandons time-varying adjustments altogether. It fixes a set of eigenvalues calibrated over an extended historical window and applies them uniformly to all subsequent covariance estimates. This static correction ignores short-term noise but captures the enduring structure of the market. Empirical tests show that AO consistently outperforms both DCC-based and other NLS techniques, suggesting that, under strong non-stationarity, long-term averages provide more reliable covariance estimates than complex dynamic models [27, 29].
@@
The AO method effectively serves as a zero-order baseline, applying fixed eigenvalue corrections without any attempt to anticipate future shifts. An ideal covariance estimator, by contrast, would move beyond denoising past returns and actively predict forthcoming covariance dynamics. NNs naturally suggest themselves for this role, since they can model highly nonlinear relationships between input data and desired outputs. Two main strategies can be used when training an NN for GMV optimization. The first is based on learning a covariance matrix estimator, which is then used to compute the portfolio weights with a Quadratic Programming (QP) minimization. This approach has the advantage of being versatile and adaptable to different minimization problems, especially if they are convex and optimally solvable. However, it has the drawback of requiring the estimation of a covariance matrix, which needs to be positive definite. In addition, the loss function should be taken as some type of distance between the predicted covariance matrix and the population covariance matrix, usually the Frobenius distance. This presents a challenge, as the Frobenius distance is not a good metric for covariance matrices in practical portfolio optimization problems [30, 31]. The second approach is to train an NN to predict portfolio weights directly. This method might allow for a straightforward optimization of the desired loss function if the NN architecture is designed in a tailored way. Depending on the architecture, the NN can internally store a representation of the covariance matrix or some simplified version. In this paper, we adopt the latter approach.
@@
At the heart of our architecture lies the correlation-denoising module (Sec. 3.2), which implements a RIE of the correlation matrix. By applying a RIE, one forces all permutational symmetries already encoded in its mathematical formulation. To perform this eigenvalue filtering efficiently, the symmetry of the eigenvalue transformation itself must be respected. We therefore take inspiration from the well-known mapping of sample eigenvalues to a one-dimensional Coulomb gas [32], in which each eigenvalue behaves as repelling identical charges confined by the population spectrum. In our implementation, the ordered sequence of sample eigenvalues is processed through a bidirectional LSTM that enforces permutation equivariance and captures the local interactions among neighboring “charges”. The network outputs a $q$-dependent NLS function that adjusts each sample eigenvalue while leaving the empirical eigenvectors intact. Two auxiliary modules prepare the raw return data for this denoising step. A learnable lag-transformation network weights look-back lags and applies soft clipping to control temporal influence and outliers (Sec. 3.1). A marginal-volatility network converts each asset’s sample standard deviation into an inverse volatility scale (Sec. 3.3). These modules work in concert to present the LSTM with input representations whose temporal dependencies and scales are optimally calibrated. Finally, the cleaned eigenvalues recombine with the inverse marginal volatilities to form an estimated inverse covariance matrix, from which GMV portfolio weights follow in closed form. Calibrating all modules jointly to minimize realized OOS variance enables the model to predict future covariance dynamics under non-stationary market regimes. This end-to-end design thus overcomes the limits of Frobenius-norm–based shrinkage and static filters and delivers a dynamically adaptive estimator of portfolio risk.
@@ 1.1 Literature review
The literature on NN portfolio optimization spans three methodological strands: predict-then-optimise pipelines, end-to-end differentiable architectures, and Reinforcement-Learning (RL) agents. We review some representative studies from every strand, briefly highlighting some design choices.
@@
End-to-end differentiable frameworks embed the optimiser within the network so that forecasting and allocation are learned jointly. Butler and Kwon,[36] place the Markowitz module inside the computational graph via implicit differentiation of the KKT system, achieving lower out-of-sample costs than those obtained with ordinary least squares. The model is trained on a 24-contract futures universe by tuning a single exponential moving-average covariance window and with a fixed risk-aversion coefficient. Zhang et al.,[37] train an LSTM to map price histories of four index ETFs directly to long-only weights, maximising the Sharpe ratio. The method withstands the early-2020 Covid-19 shock. The set of assets comprises a relatively small ETF set. Costa and Iyengar,[38] embed a distributionally robust mean–variance layer that back-propagates through investor-risk and ambiguity-set parameters, outperforming equal-weight and classical pipelines on 2000–2021 US equities. The sample spans 20 surviving S&P 500 stocks. Uysal et al.,[39] integrate a differentiable risk-budgeting layer and raise the Sharpe ratio from 0.79 for nominal risk-parity to 1.24 for a portfolio of seven US ETFs (2011–2021).

### v3

@@ Abstract
We develop a rotation-invariant neural network that provides the global minimum-variance portfolio by jointly learning how to lag-transform historical returns and marginal volatilities and how to regularise the eigenvalues of large equity covariance matrices. This explicit mathematical mapping offers clear interpretability of each module’s role, so the model cannot be regarded as a pure black box. The architecture mirrors the analytical form of the global minimum-variance solution yet remains agnostic to dimension, so a single model can be calibrated on panels of a few hundred stocks and applied, without retraining, to one thousand US equities, a cross-sectional jump that indicates robust generalization capability. The loss function is the future short-term realized minimum variance and is optimized end-to-end on real returns. In out-of-sample tests from January 2000 to December 2024, the estimator delivers systematically lower realized volatility, smaller maximum drawdowns, and higher Sharpe ratios than the best competitors, including state-of-the-art non-linear shrinkage, and these advantages persist across both short and long evaluation horizons despite the model’s training focus is short-term. Furthermore, although the model is trained end-to-end to produce an unconstrained minimum-variance portfolio, we show that its learned covariance representation can be used in general optimizers under long-only constraints with virtually no loss in its performance advantage over competing estimators. These advantages persist when the strategy is executed under a highly realistic implementation framework that models market orders at the auctions, empirical slippage, exchange fees, and financing charges for leverage, and they remain stable during episodes of acute market stress.
@@ 1 Introduction
Portfolio management has been fundamental to financial practice since the earliest securities markets emerged. Following Markowitz’s introduction of Modern Portfolio Theory (MPT) [60], academic and industry efforts have prioritized the construction of portfolios that optimize risk‐adjusted returns. This framework employs historical return data and estimates of pairwise asset dependencies, typically captured by covariance matrices, to determine asset weightings.
@@
The literature on machine learning–based portfolio optimization is vast and rapidly expanding [55, 83]; a comprehensive review is provided in Sec. 1.1, and here we summarize only high-level themes. Existing studies can be grouped into three main categories: span predict-then-optimize pipelines [61], end-to-end differentiable frameworks [14, 16, 75], and Reinforcement-Learning (RL) agents [15, 40, 79]. Overall, machine learning–based portfolio optimization is highly challenging with several limitations still to be overcome. In fact, some studies employ small or static asset universes that limit evidence of scalability and robustness [61, 83, 34, 75]. Benchmarks are in some cases infrequent or sub-optimal, making it difficult to compare performance improvements across studies [55]. Moreover, empirical protocols can be sensitive to data leakage, for instance through survivorship-restricted investable universes [34, 16, 80] or, more generally, through look-ahead bias in feature construction and evaluation. These challenges constitute the principal obstacles to the practical deployment of machine learning in portfolio optimization.
Here we propose an end-to-end neural network for Global Minimum-Variance (GMV) portfolio optimization, designed to operate in an asset-selection setting that closely matches professional investment practice. The approach is dimensionally agnostic: a model calibrated on a few hundred stocks can be deployed on substantially larger universes without retraining. We evaluate performance using a realistic backtesting protocol that reproduces the operational life cycle of a cash-and-margin account, including long-only constraints, transaction costs, and other trading frictions. In Out-Of-Sample (OOS) tests on U.S. equities (2000–2024), the resulting portfolios achieve lower realized volatility, smaller drawdowns, and higher Sharpe ratios than widely used state-of-the-art covariance estimators, while scaling reliably from a few hundred to approximately 1,000 assets.
Methodologically, we recast the classical GMV solution as a differentiable rotation-invariant pipeline and implement its key steps as a modular neural architecture. Returns for $n$ assets over $\Delta t_{\textrm{in}}$ observations are first processed by a learnable lag-transformation with soft clipping, after which a multivariate denoising block outputs a cleaned correlation structure while adapting to the aspect ratio $q_{\textrm{in}}:=n/\Delta t_{\textrm{in}}$, in parallel, a feed-forward module estimates inverse marginal volatilities. All components are trained jointly using the realized OOS variance of the resulting GMV portfolio as the learning signal, so that preprocessing, volatility scaling, and eigenvalue regularization are optimized end-to-end rather than selected a priori. The main architectural challenge lies in the multivariate denoising stage, where symmetry and validity constraints inherent to correlation matrices severely restrict admissible parameterizations. By embedding these constraints through a rotation-invariant design and exploiting Random Matrix Theory (RMT)-informed spectral regularities, the learned parameters can leverage known symmetries and dependencies, facilitating transfer across asset universes of different dimensionality.
@@
A key aspect of our model is associated with the determination of the covariance matrix. When estimated from finite samples, empirical covariance matrices tend to be ill‐conditioned and sample covariance presents estimation errors [69, 43, 48], often resulting in suboptimal portfolio allocations. A widely adopted remedy is to apply Rotation Invariant Estimators (RIEs), which adjust the sample eigenvalues while retaining the sample eigenvectors [13]. Within this framework, optimality is measured by the Mean Squared Error (MSE), equivalent to the Frobenius norm distance between the estimator and the true covariance matrix. The theoretical optimum is the so‐called Oracle Estimator (OE), defined as the covariance matrix that minimizes this Frobenius distance. In practice, however, the OE cannot be realized, since it requires prior knowledge of the unknown true covariance.
@@
One of the earliest covariance regularization techniques is eigenvalue clipping. Laloux et al. [43] showed that the bulk of the eigenvalue spectrum of empirical financial covariance matrices closely follows the Marchenko–Pastur distribution for random noise. Eigenvalue clipping therefore replaces all sample eigenvalues that lie within this noise bulk by their average value. Later, Ledoit and Wolf [48] introduced the Linear Shrinkage (LS) estimator, which constructs a convex combination of the sample covariance matrix and a diagonal target matrix. This estimator remains well–defined when the number of assets $n$ exceeds the number of observations $\Delta t_{\textrm{in}}$. Its optimality in minimizing MSE holds strictly under the assumption that the true covariance matrix is drawn from an inverse–Wishart distribution [46]. In the last decade, Ledoit and collaborators introduced the NonLinear Shrinkage (NLS) estimator as an optimal method for covariance regularization [49]. Deriving the analytical form of the Ledoit–Péché (LP) estimator is challenging because it requires detailed knowledge of the sample eigenvector distribution [47]. The Quantized Eigenvalues Sampling Transform (QuEST) delivers the most accurate numerical implementation of the LP estimator by discretizing the limiting spectral distribution [50], but it incurs substantial computational cost. To alleviate this burden, the direct kernel estimator [51] and the inverse–Wishart regularization approach [13] offer closed-form approximations that closely mimic the LP solution. More recently, Ledoit and Wolf proposed the Quadratic-Inverse Shrinkage (QIS) estimator [53], which is considered the state-of-the-art technique in financial covariance filtering by combining computational efficiency with high estimation accuracy.
@@
Although the theoretical framework for covariance filtering is well established, its practical implementation remains challenging due to the pronounced non-stationarity of financial covariances. Market regimes evolve continuously under shifting economic conditions and investor behavior, so a covariance matrix estimated from past returns may poorly reflect current asset dependencies. Moreover, NLS theory assumes that sampling noise is the sole source of discrepancy between the sample and true covariance matrices and relies on the thermodynamic limit in which the number of assets $n$ and observations $\Delta t_{\textrm{in}}$ both tend to infinity with fixed aspect ratio $q$. In practice, extending $\Delta t_{\textrm{in}}$ to satisfy this limit only exacerbates non-stationarity, as longer historical windows encompass multiple structural shifts. Finally, the optimality of the LP estimator depends on the existence of finite twelfth-order moments of asset returns [47], yet empirical evidence indicates that equity returns exhibit heavy tails with tail exponents near four, rendering higher-order moments effectively unbounded. Methods based on $k$-fold Cross‐Validation (CV) have been proposed to relax the stringent assumptions of nonlinear shrinkage theory [3, 45, 44, 46, 65]. These CV estimators achieve accuracy comparable to more complex schemes while permitting straightforward adaptations to specific data features. For example, Tan et al. [71] introduced an exponential decay factor into the return series within the CV framework. By applying exponential weights, one can maintain a long calibration window, approaching the thermodynamic limit, while emphasizing the most recent observations. This modification partially resolves the bias-variance trade‐off caused by non‐stationarity and structural shifts in financial covariances.
@@
An alternative to RIEs employs hierarchical clustering to filter the empirical correlation matrix [73, 7, 8]. Unlike RIEs, clustering‐based filters modify both eigenvalues and eigenvectors by aggregating assets into empirically determined groups. This approach exploits the fact that financial assets are often organized into a clear hierarchy of sectors and sub‐sectors [58]. Although it lacks full theoretical support, clustering filters have, in some instances, yielded superior out‐of‐sample portfolio performance compared with RIEs [72, 63, 7, 8, 30]. By enforcing the observed hierarchical relationships, these methods produce the most parsimonious covariance approximation that exactly reproduces the data’s cluster structure.
@@
The Engle Dynamical Conditional Covariance (DCC) model [23] represents a principled attempt to accommodate non‐stationarity by modeling time‐varying correlations through Generalized AutoRegressive Conditional Heteroscedasticity (GARCH)‐type dynamics. It specifies separate equations for the conditional variances and then characterizes the evolving correlations of the standardized returns via a time-varying correlation matrix. In principle, this approach should adapt to shifting market regimes and capture persistence in asset comovements. In practice, however, DCC is computationally intensive, often suffers from convergence issues, and yields OOS portfolios whose risk–return profiles rarely surpass those obtained with simpler static filters [10]. In contrast, the Average Oracle (AO) method [6, 10] abandons time‐varying adjustments altogether. It fixes a set of eigenvalues calibrated over an extended historical window and applies them uniformly to all subsequent covariance estimates. This static correction ignores short‐term noise but captures the enduring structure of the market. Empirical tests show that AO consistently outperforms both DCC‐based and other NLS techniques, suggesting that, under strong non‐stationarity, long‐term averages provide more reliable covariance estimates than complex dynamic models [10, 26].
@@
The AO method effectively serves as a zero‐order baseline, applying fixed eigenvalue corrections without any attempt to anticipate future shifts. An ideal covariance estimator, by contrast, would move beyond denoising past returns and actively predict forthcoming covariance dynamics. NNs naturally suggest themselves for this role, since they can model highly nonlinear relationships between input data and desired outputs. Two main strategies can be used when training an NN for GMV optimization. The first is based on learning a covariance matrix estimator, which is then used to compute the portfolio weights with a Quadratic Programming (QP) minimization. This approach has the advantage of being versatile and adaptable to different minimization problems, especially if they are convex and optimally solvable. However, it has the drawback of requiring the estimation of a covariance matrix, which needs to be positive definite. In addition, the loss function should be taken as some type of distance between the predicted covariance matrix and the population covariance matrix, usually the Frobenius distance. This presents a challenge, as the Frobenius distance is not a good metric for covariance matrices in practical portfolio optimization problems [11, 9]. The second approach is to train an NN to predict portfolio weights directly. This method might allow for a straightforward optimization of the desired loss function if the NN architecture is designed in a tailored way. Depending on the architecture, the NN can internally store a representation of the covariance matrix or some simplified version. In this paper, we adopt the latter approach.
@@
At the heart of our architecture lies the correlation‐denoising module (Sec. 3.2), which implements a RIE of the correlation matrix. By applying a RIE, one forces all permutational symmetries already encoded in its mathematical formulation. To perform this eigenvalue filtering efficiently, the symmetry of the eigenvalue transformation itself must be respected. We therefore take inspiration from the well‐known mapping of sample eigenvalues to a one‐dimensional Coulomb gas [28], in which each eigenvalue behaves as repelling identical charges confined by the population spectrum. In our implementation, the ordered sequence of sample eigenvalues is processed through a bidirectional LSTM that enforces permutation equivariance and captures the local interactions among neighboring “charges”. The network outputs a $q$-dependent NLS function that adjusts each sample eigenvalue while leaving the empirical eigenvectors intact. Two auxiliary modules prepare the raw return data for this denoising step. A learnable lag‐transformation network weights look-back lags and applies soft clipping to control temporal influence and outliers (Sec. 3.1). A marginal‐volatility network converts each asset’s sample standard deviation into an inverse volatility scale (Sec. 3.3). These modules work in concert to present the LSTM with input representations whose temporal dependencies and scales are optimally calibrated. Finally, the cleaned eigenvalues recombine with the inverse marginal volatilities to form an estimated inverse covariance matrix, from which GMV portfolio weights follow in closed form. Calibrating all modules jointly to minimize realized OOS variance enables the model to predict future covariance dynamics under non‐stationary market regimes. This end‐to‐end design thus overcomes the limits of Frobenius‐norm–based shrinkage and static filters and delivers a dynamically adaptive estimator of portfolio risk.
@@ 1.1 Literature Review
The literature on NN portfolio optimization spans three methodological strands: predict-then-optimise pipelines, end-to-end differentiable architectures, and RL agents. We review some representative studies from every strand, briefly highlighting some design choices.
@@
End-to-end differentiable frameworks embed the optimiser within the network so that forecasting and allocation are learned jointly. Butler and Kwon,[14] place the Markowitz module inside the computational graph via implicit differentiation of the KKT system, achieving lower OOS costs than those obtained with ordinary least squares. The model is trained on a 24-contract futures universe by tuning a single exponential moving-average covariance window and with a fixed risk-aversion coefficient. Zhang et al.,[83] train an LSTM to map price histories of four index ETFs directly to long-only weights, maximising the Sharpe ratio. The method withstands the early-2020 Covid-19 shock. The set of assets comprises a relatively small ETF set. Costa and Iyengar,[16] embed a distributionally robust mean–variance layer that back-propagates through investor-risk and ambiguity-set parameters, outperforming equal-weight and classical pipelines on 2000–2021 US equities. The sample spans 20 surviving S&P 500 stocks. Uysal et al.,[75] integrate a differentiable risk-budgeting layer and raise the Sharpe ratio from 0.79 for nominal risk-parity to 1.24 for a portfolio of seven US ETFs (2011–2021).